OpenStudy (maeganr):

What is the value of -6a^-3b^2 for a=-2 and b=4?

OpenStudy (anonymous):

wherever you see an a put -2 and wherever you see a b put a 4

OpenStudy (maeganr):

I know that. I need to know how to solve that then.

OpenStudy (anonymous):

well start putting it in and upload what you have.

OpenStudy (maeganr):

-6(-2)^-3(4)^2

OpenStudy (anonymous):

ok so in other words you have \[\frac{-6}{2^{-3}}*4^2\]

OpenStudy (anonymous):

gotta go but if you expand and simplify then you can solve it.

OpenStudy (maeganr):

Oh. Kay?

OpenStudy (maeganr):

@kelliegirl33 @Hope_nicole @tkhunny

OpenStudy (anonymous):

ok, i'm back (-2)^-3 = -1/8 4^2 = 16 (-6)(-1/8)(16) -6*-2 12

OpenStudy (anonymous):

still need help?

OpenStudy (maeganr):

Lil bit. @Hope_nicole

OpenStudy (maeganr):

Nevermind